# Frostman random variables, entropy inequalities, and applications

### v3

@@ 2.1 Notations
Let $(\Omega,\mathfrak{B},P)$ be a probability space. A random variable $X$ is a measurable function from $(\Omega,\mathfrak{B},P)$ to another measuable space $(\Omega^{\prime},\mathfrak{B}^{\prime})$, on which a pushforward probability measure $\mu$ is induced, i.e., $\mu(B^{\prime})=P(X\in B^{\prime})$ for any $B^{\prime}\in\mathfrak{B}^{\prime}$. We will say that $\mu$ is the associated distribution measure of $X$. We denote the range of $X$ by $range(X)$, which collects all the possible values that $X$ can take. In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, the Euclidean norm $|\cdot|$, and the Lebesgue measure $\mathcal{L}^{m}(\cdot)$. The dimension $m$ is always regarded as an absolute constant. The range of any random variable is assumed to be a bounded set. When two random variables $X$ and $Y$ have the same distribution, we write $X\equiv Y$.
@@
For a measure $\mu$ and a Borel set $E$ with $\mu(E)>0$, we define the conditional measure of $\mu$ with respect to $E$ by $\mu_{E}(B)=\frac{\mu(B\cap E)}{\mu(E)}$ for any Borel set $B$. In particular, when $\mu=P$, the conditional probability measure given an event $E$ is denoted by $P_{E}$. For a random varaible $X$ and such an $E$, we denote by $X_{E}$ the random variable satisfying $X_{E}\equiv(X|E)$ (see Lemma A.2). That is to say,
@@
For a measurable function $f:\,(\Omega^{\prime},\mathcal{B}^{\prime})\rightarrow(\Omega^{\prime\prime},\mathfrak{B}^{\prime\prime})$ between measurable spaces, the random variable $f(X)$ is defined to be the composition of $f$ and $X$. The domain and range of $f$ are denoted by $domain(f)$ and $range(f)$, respectively. Throughout this paper, all the functions are assumed to be measurable. And whenever we write $f(X)$ with $X$ a random variable and $f$ a function, we always mean that $domain(f)\supseteq range_{+}(X)$, where the values taken on a zero-measure set is neglectable.
@@ 2.2 Entropy and basic properties
In above definition, the convention $0\log 0=0$ is used for convenience. We also say $\mu(B)H(\mu_{B},\mathcal{A})=0$ when $\mu(B)=0$, although $\mu_{B}$ is not defined in this situation. Intuitively, the entropy of a random variable will tell us how hard it is to determine a value. Before stating the first lemma, we introduce the following notation:
@@
Let $X$ and $Y$ be discrete random variables, and $E$ be an event. The entropy of $X$ is defined as
@@
So, by applying Lemma 2.4 (iii) and (8), we obtain
@@
Let us consider the the maximum value of the function
@@
That is to say, when the other k1k-1 variables are given, the small $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take. Now one can verify that the maximum value of $f$ is taken at $(p_{1},\ldots,p_{k+1})$ with
@@ 2.3 Discretized entropy and properties
For each $n\in\mathbb{N}$ and $m\in\mathbb{N}$, let us define a function $D_{n,m}:\,{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}^{m}$ as the following. For any $(x_{1},\ldots,x_{m})\in{\mathbb{R}}^{m}$, if $k_{i}$ is the unique integer such that $\frac{k_{i}}{2^{n}}\leq x_{i}<\frac{k_{i}+1}{2^{n}}$ $(i=1,\ldots,m)$, then we define $D_{n,m}(x_{1},\ldots,x_{m})=(\frac{k_{1}}{2^{n}},\ldots,\frac{k_{m}}{2^{n}})$. That is to say, $D_{n,m}(x)$ locates the left endpoints of the dyadic intervals that the $x_{i}$’s fall in. For convenience, we will simplify $D_{n,m}$ as $D_{n}$, where the dimension $m$ is determined from the context. For any random variable $X$, the expression $D_{n}(X)$ is a discrete random variable that takes values in the set of points whose components are all dyadic rationals. Moreover, if $X_{1},\ldots,X_{k}$ are independent (or identically distributed) random variables, then $D_{n}(X_{1}),\ldots,D_{n}(X_{k})$ are also independent (or identically distributed, respectively).
@@
Moreover, the discretized entropy only depends on the discretised distribution. That is to say, if $D_{n}(X)$ and $D_{n}(Y)$ are identically distributed, then $H_{n}(X)=H_{n}(Y)$.
@@
Let $\kappa$ be a positive real number, and $f$ be a function.
(i) We say that $f$ is $\kappa$-Lipschitz if $|f(x)-f(y)|\leq\kappa|x-y|$ for any $x,y\in domain(f)$.
(ii) We say that $f$ is $\kappa$-bi-Lipschitz if $f$ is invertible, and both $f$ and $f^{-1}$ are $\kappa$-Lipschitz.
(iii) If $f$ is $\kappa$-Lipschitz, and there exists a decomposition of $domain(f)=\mathcal{R}_{1}\cup\ldots\cup\mathcal{R}_{l}$, where $l$ is a positive integer viewed as an absolute constant, such that each $f|_{\mathcal{R}_{i}}$ has an $\kappa$-Lipschitz inverse $(i=1,\ldots,l)$, then we say that $f$ is weakly $\kappa$-bi-Lipschitz.
@@
Let $D$ be an open convex region in ${\mathbb{R}}^{m}$. Let $f:\,D\rightarrow{\mathbb{R}}$ be a function such that $|\frac{\partial f}{\partial x_{i}}(x_{1},\ldots,x_{m})|\leq M$ for all $(x_{1},\ldots,x_{m})\in D$. Then, for any $(x_{1},\ldots,x_{m})$ and $(y_{1},\ldots,y_{m})$ in $D$, there is some $(\xi_{1},\ldots,\xi_{m})\in D$ such that
@@
Let $\alpha,c_{1},c_{2}$ be real numbers with $\alpha\neq 0$ and $0<c_{2}<c_{1}$. Let $f(x)=\alpha x^{2}$ with domain $c_{1},-c_{2}]\cup[c_{2},c_{1}]$. Note that, by the mean value theorem,
@@
whenever $x,y\in[-c_{1},-c_{2}]$, or $x,y\in[c_{2},c_{1}]$. One asserts that $f$ is weakly $\max\left\{(2|\alpha|c_{2})^{-1},\,2|\alpha|c_{1}\right\}$-bi-Lipschitz.
@@
Let $X$, $Y$ be random variables and $\kappa$ be a postive real number. We say that $Y$ is a $\kappa$-Lipschitz function (or $\kappa$-bi-Lipschitz function, or weakly $\kappa$-bi-Lipschitz function) of $X$ if $Y=f(X)$ for some $\kappa$-Lipschitz function (or $\kappa$-bi-Lipschitz function, or weakly $\kappa$-bi-Lipschitz function, respectively) $f$ with $domain(f)\supseteq range_{+}(X)$.
@@
(i) If $f$ is $\kappa$-Lipschitz, then for all $n\in\mathbb{N}$,
@@
(ii) If $g$ be weakly $\kappa$-bi-Lipschitz, then for all $n\in\mathbb{N}$,
$H_{n}(Y|g(Y)))\ll\log\kappa.$
The above implied constants are independent of $n$, and the concrete expressions of the functions and random variables.
@@
Assume that $domain(g)\subseteq{\mathbb{R}}^{m}$ and $range(g)\subseteq{\mathbb{R}}^{m^{\prime}}$, where $m,m^{\prime}$ are viewed as absolute constants. When $D_{n}(g(Y))$ is determined, one knows that $g(Y)$ takes value in a ball with radius $O_{m^{\prime}}(2^{-n})$. By the definition of weakly $\kappa$-bi-Lipschitz functions, we can locate $Y$ to be lying in the pre-image of this ball, which is a union of at most $O(1)$ balls with radius $O_{m^{\prime}}(\kappa\cdot 2^{-n})$. So $D_{n}(Y)$ has at most $O_{m,m^{\prime}}(\kappa^{m})$ choices. It follows from Lemma 2.4(ii) that
$H_{n}(Y|g(Y))\leq\log(O_{m,m^{\prime}}(\kappa^{m}))\ll\log\kappa.$
@@
(i) Let $X$ and $Y$ be random variables such that $X$ is an $\kappa$-Lipschitz function of $Y$. Then, for all $n\in\mathbb{N}$,
@@
(ii) Let $X$ and $Y$ be random variables such that $X$ is a weakly $\kappa$-bi-Lipschitz function of $Y$. Then, for all $n\in\mathbb{N}$,
$H_{n}(X)=H_{n}(Y)+O(\log\kappa).$
@@
Let $X$ be a random variable with $range_{+}(X)\subseteq[-c_{1},-c_{2}]\cup[c_{2},c_{1}]$, where $0<c_{2}<c_{1}$. Let $\alpha\in{\mathbb{R}}^{\ast}$. One can deduce from Example 2.12 and Lemma 2.15 that
$H_{n}(\alpha X^{2})=H_{n}(X)+O\big{(}\max\{1,\log c_{1},\log(1/c_{2})\}\big{)}+O(\big{|}\log|\alpha|\big{|}).$
Example 2.17.
Let $m\geq 1$. Suppose that $A$ is an $m\times m$ real matrix and $X,X_{1},X_{2}$ are multivariate real random variables of dimension $m$. Note that $|A(X_{1}-X_{2})|\leq\|A\|\cdot|X_{1}-X_{2}|$, where $\|A\|$ is the operator norm of $A$. Lemma 2.15 shows that $H_{n}(AX)\leq H_{n}(X)+O_{A}(1)$. In particular, if $A$ is invertible, then $H_{n}(AX)=H_{n}(X)+O_{A}(1)$.
@@
Let $X$, $Y$ be random variables, and $f$ be a $\kappa$-Lipschitz function with $range(f)\subseteq{\mathbb{R}}^{m}$, where $m$ is viewed as an absolute constant. Suppose that $|X-Y|\leq K$ with probability $1$. Then, for all $n\in\mathbb{N}$ with $2^{-n}\leq K$,
@@
$\max\{H_{n}(X),H_{n}(Y)\}\leq H_{n}(X+Y)+O(1)$ (13)
for all $n\in\mathbb{N}$, provided that $X$ and $Y$ are independent. In the rest of this paper, we may apply Lemma 2.20 without mention it.
In [16], Ruzsa studied some entropy inequalities that involving sums of random variables. In the following, more inequalities for discretized entropy are presented.
@@
Let $X,Y$ and $Z$ be independent real random variables. Then, for all $n\in\mathbb{N}$,
@@
Without independence, the follow lemma can be used as an alternative.
@@
Let $k,l\in\mathbb{Q}^{\ast}$, and $X$ and $Y$ be real random variables that are not necessarily independent. Then there exist integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
@@
Let $k\in\mathbb{Z}^{\ast}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k)>0$ and $d_{2}=d_{2}(k)\geq 0$ such that
@@
For $k=1$, the conclusion holds for $d_{1}=1$ and $d_{2}=0$. For k=1k=-1, Lemma 2.24 shows that we can take $d_{1}=3$ and $d_{2}=5$. Now we apply induction on $|k|$. Assume that the conclusion has been proved for $k$ with $|k|=m$, where $m\geq 1$. We consider the situation $k=\pm(m+1)$.
@@
Let $k,l\in\mathbb{Z}^{\ast}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}$ $\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\mathcal{I}_{n}(X,Y)+O(1).$
@@
Applying Lemma 2.26 with the random variable $(X,lY)$, we obtain that
@@
$\qquad\leq d_{1}(k,1)\left(H_{n}(X+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}(k,1)\mathcal{I}_{n}(X,Y)+O_{k,l}(1).$
Applying Lemma 2.26 again, with the random variable $(Y,X)$, we find that
@@
Since $k,l\in\mathbb{Q}^{\ast}$, we can write $k=k^{\prime}/t$ and $l=l^{\prime}/t$ for some $k^{\prime},l^{\prime},t\in\mathbb{Z}^{\ast}$. It satisfies that $H_{n}(kX+lY)=H_{n}(k^{\prime}X+l^{\prime}Y)+O_{t}(1)$. The conclusion follows from Lemma 2.27. ∎
@@
for any $n\in\mathbb{N}$. Here, the case that $k=0$ or $l=0$ can be proved by using (13).
@@
Let $X,Y$ be independent random variables, and $f,g$ be weakly $\kappa$-bi-Lipschitz real-valued functions. Suppose that $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$. Then
$H_{n}(f(X_{1})+f(X_{2})+g(Y_{1})+g(Y_{2}))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa),$
$H_{n}(f(X_{1})-f(X_{2})+g(Y_{1})-g(Y_{2}))$ $\leq 7H_{n}(f(X)+g(Y))-3H_{n}(X)-3H_{n}(Y)+O(\log\kappa),$
$H_{n}(f(X_{1})+f(X_{2})-g(Y_{1})-g(Y_{2}))$ $\leq 9H_{n}(f(X)+g(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\kappa).$

### v4

@@ 2.1 Notations
Let $(\Omega,\mathfrak{B},P)$ be a probability space. A random variable $X$ is a measurable function from $(\Omega,\mathfrak{B},P)$ to another measurable space $(\Omega^{\prime},\mathfrak{B}^{\prime})$, on which a pushforward probability measure $\mu$ is induced, i.e., $\mu(B^{\prime})=P(X\in B^{\prime})$ for any $B^{\prime}\in\mathfrak{B}^{\prime}$. We will say that $\mu$ is the associated distribution measure of $X$. We denote the range of $X$ by $range(X)$, which collects all the possible values that $X$ can take. In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, the Euclidean norm $|\cdot|$, and the Lebesgue measure $\mathcal{L}^{m}(\cdot)$, where the dimension $m$ is considered as an absolute constant. The range of any random variable is assumed to be a bounded set. When two random variables $X$ and $Y$ have the same distribution, we write $X\equiv Y$.
@@
For a measure $\mu$ and a Borel set $E$ with $\mu(E)>0$, we define the conditional measure of $\mu$ with respect to $E$ by $\mu_{E}(B)=\frac{\mu(B\cap E)}{\mu(E)}$ for any Borel set $B$. In particular, when $\mu=P$, the conditional probability measure given an event $E$ is denoted by $P_{E}$. For a random variable $X$ and such an event $E$, we denote by $X_{E}$ a random variable satisfying $X_{E}\equiv(X|E)$ (see Lemma A.2). That is to say,
@@
For a measurable function $f:\,(\Omega^{\prime},\mathcal{B}^{\prime})\rightarrow(\Omega^{\prime\prime},\mathfrak{B}^{\prime\prime})$ between measurable spaces, the random variable $f(X)$ is defined to be the composition of $f$ and $X$. The domain and range of $f$ are denoted by $domain(f)$ and $range(f)$, respectively. Throughout this paper, all the functions are assumed to be measurable. And whenever we write $f(X)$ with $X$ a random variable and $f$ a function, we always mean that $domain(f)\supseteq range_{+}(X)$, where the values taken on a zero-measure set is negligible.
@@ 2.2 Entropy and basic properties
In this definition, we use the convention $0\log 0=0$. We also interpret $\mu(B)H(\mu_{B},\mathcal{A})$ as 0 when $\mu(B)=0$, even though $\mu_{B}$ is not defined in this situation. Intuitively, the entropy of a random variable will tell us how hard it is to determine a value. Before stating the first lemma, we introduce the following notation:
@@
Let $X$ and $Y$ be discrete random variables, and $E$ be an event with positive probability. The entropy of $X$ is defined as
@@
Hence, by applying Lemma 2.4 (iii) and (5), we obtain
@@
Let us analyze the the maximum value of the function
@@
In other words, when the other k1k-1 variables are given, the small $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take. Now we can verify that the maximum value of $f$ is taken at $(p_{1},\ldots,p_{k+1})$ with
@@ 2.3 Discretized entropy and properties
For each $n\in\mathbb{N}$ and $m\in\mathbb{N}$, define $D_{n,m}:\,{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}^{m}$ by
$D_{n,m}(x_{1},\ldots,x_{m})=\left(\frac{k_{1}}{2^{n}},\ldots,\frac{k_{m}}{2^{n}}\right),$
where for each $i$ the integer $k_{i}$ is uniquely determined by $\frac{k_{i}}{2^{n}}\leq x_{i}<\frac{k_{i}+1}{2^{n}}$. So, $D_{n,m}(x)$ locates the left endpoints of the dyadic intervals that the $x_{i}$’s fall in. For convenience, we often write $D_{n}$ instead of $D_{n,m}$ when the dimension is clear from the context. For any random variable $X$, the expression $D_{n}(X)$ is a discrete random variable that takes values in the set of points whose components are all dyadic rationals. Moreover, if $X_{1},\ldots,X_{k}$ are independent (or identically distributed) random variables, then $D_{n}(X_{1}),\ldots,D_{n}(X_{k})$ are independent (or identically distributed, respectively).
@@
Moreover, the discretized entropy only depends on the discretized distribution. If $D_{n}(X)$ and $D_{n}(Y)$ are identically distributed, then $H_{n}(X)=H_{n}(Y)$.
@@
Let $\kappa$ be a positive real number, $f$ be a function, and $\mathcal{R}$ be a region contained in $domain(f)$.
(i) We say that $f$ is $\kappa$-Lipschitz on $\mathcal{R}$ if $|f(x)-f(y)|\leq\kappa|x-y|$ for any $x,y\in\mathcal{R}$.
(ii) We say that $f$ is $\kappa$-bi-Lipschitz on $\mathcal{R}$ if $f|_{\mathcal{R}}$ is invertible, with $f$ and $(f|_{\mathcal{R}})^{-1}$ being $\kappa$-Lipschitz on $\mathcal{R}$ and $f(\mathcal{R})$, respectively.
(iii) If $f$ is $\kappa$-Lipschitz on $\mathcal{R}$, and there exists a decomposition $\mathcal{R}=\mathcal{R}_{1}\cup\ldots\cup\mathcal{R}_{d}$, where $d$ is a positive integer, such that each $f|_{\mathcal{R}_{i}}$ has an $\kappa$-Lipschitz inverse on $f(\mathcal{R}_{i})$ $(i=1,\ldots,d)$, then we say that $f$ is weakly $(\kappa,d)$-bi-Lipschitz.
For simplicity, we will abbreviate weakly $(\kappa;O(1))$-bi-Lipschitz as weakly $\kappa$-bi-Lipschitz. And we may omit the phrase “on $\mathcal{R}$” if it is clear from the context.
@@
Let $\mathcal{R}$ be an open convex region in ${\mathbb{R}}^{m}$. Let $f:\,\mathcal{R}\rightarrow{\mathbb{R}}$ be a function such that $|\frac{\partial f}{\partial x_{i}}(x_{1},\ldots,x_{m})|\leq M$ for all $(x_{1},\ldots,x_{m})\in\mathcal{R}$. Then, for any $(x_{1},\ldots,x_{m})$ and $(y_{1},\ldots,y_{m})$ in $\mathcal{R}$, there is some $(\xi_{1},\ldots,\xi_{m})\in\mathcal{R}$ such that
@@
Let $\alpha,c_{1},c_{2}$ be real numbers with $\alpha\neq 0$ and $0<c_{2}<c_{1}$. Let $f(x)=\alpha x^{2}$. Note that, by the mean value theorem,
@@
whenever $x,y\in[-c_{1},-c_{2}]$, or $x,y\in[c_{2},c_{1}]$. One asserts that $f$ is weakly $\max\left\{(2|\alpha|c_{2})^{-1},\,2|\alpha|c_{1}\right\}$-bi-Lipschitz on $c_{1},-c_{2}]\cup[c_{2},c_{1}]$.
The following lemma is a generalization of above example, which follows immediately by mean value theorem.
Lemma 2.13.
Let $\mathcal{R}\subseteq{\mathbb{R}}$ and $f:\,\mathcal{R}\rightarrow{\mathbb{R}}$ be a differentiable function. Suppose that $\mathcal{R}$ is a union of $d$ intervals such that $f$ is monotonic on each of the intervals. Let $M=\sup\nolimits_{x\in\mathcal{R}}|f^{\prime}(x)|$ and $m=\inf\nolimits_{x\in\mathcal{R}}|f^{\prime}(x)|$. Then $f$ is weakly $(\max\{M,m^{-1}\},d)$-bi-Lipschitz on $\mathcal{R}$.
@@
Let $X$, $Y$ be random variables and $\kappa$ be a positive real number. We say that $Y$ is a $\kappa$-Lipschitz function (or $\kappa$-bi-Lipschitz function, or weakly $(\kappa,d)$-bi-Lipschitz function) of $X$ if $Y=f(X)$ for some function $f$ that is $\kappa$-Lipschitz (or $\kappa$-bi-Lipschitz, or weakly $(\kappa,d)$-bi-Lipschitz, respectively) on $range_{+}(X)$.
@@
(i) If $f$ is $\kappa$-Lipschitz on $range_{+}(X)$, then for all $n\in\mathbb{N}$,
@@
(ii) If $g$ is weakly $(\kappa,d)$-bi-Lipschitz on $range_{+}(Y)$, then for all $n\in\mathbb{N}$,
$H_{n}(Y|g(Y)))\ll\log\kappa+\log d.$
The implied constants are independent of $n$ and of the particular functions and random variables involved.
@@
Assume that $domain(g)\subseteq{\mathbb{R}}^{m}$ and $range(g)\subseteq{\mathbb{R}}^{m^{\prime}}$, where $m,m^{\prime}$ are viewed as absolute constants. When $D_{n}(g(Y))$ is determined, one knows that $g(Y)$ takes value in a ball with radius $O_{m^{\prime}}(2^{-n})$. By the definition of weakly $(\kappa,d)$-bi-Lipschitz functions, we can locate $Y$ to be lying in the pre-image of this ball, which is a union of at most $d$ balls with radius $O_{m^{\prime}}(\kappa\cdot 2^{-n})$. So $D_{n}(Y)$ has at most $O_{m,m^{\prime}}(d\kappa^{m})$ choices. It follows from Lemma 2.4(ii) that
$H_{n}(Y|g(Y))\leq\log(O_{m,m^{\prime}}(d\kappa^{m}))\ll\log\kappa+\log d.$
@@
(i) Let $X$ and $Y$ be random variables such that $X$ is a $\kappa$-Lipschitz function of $Y$. Then, for all $n\in\mathbb{N}$,
@@
(ii) Let $X$ and $Y$ be random variables such that $X$ is a weakly $(\kappa,d)$-bi-Lipschitz function of $Y$. Then, for all $n\in\mathbb{N}$,
$H_{n}(X)=H_{n}(Y)+O(\log\kappa)+O(\log d).$
@@
Let $m\geq 1$. Suppose that $A$ is an $m\times m$ real matrix and $X,X_{1},X_{2}$ are multivariate real random variables of dimension $m$. Note that $|A(X_{1}-X_{2})|\leq\|A\|\cdot|X_{1}-X_{2}|$, where $\|A\|$ is the operator norm of $A$. Lemma 2.16 shows that $H_{n}(AX)\leq H_{n}(X)+O_{A}(1)$. When $A$ is invertible, we further have $H_{n}(AX)=H_{n}(X)+O_{A}(1)$.
In particular, for any random variable $X$, $Y$, we have
$H_{n}(X,Y)=H_{n}(X,X+Y)+O(1)\leq H_{n}(X)+H_{n}(X+Y)+O(1).$ (8)
@@
Let $X$, $Y$ be random variables. Suppose that $f$ is $\kappa$-Lipschitz on $range_{+}(X)\cup range_{+}(Y)$, and $range(f)\subseteq{\mathbb{R}}^{m}$, where $m$ is viewed as an absolute constant. Suppose that $|X-Y|\leq K$ with probability $1$. Then, for all $n\in\mathbb{N}$ with $2^{-n}\leq K$,
@@
$\frac{H_{n}(X)+H_{n}(Y)}{2}\leq\max\{H_{n}(X),H_{n}(Y)\}\leq H_{n}(X+Y)+O(1)$ (9)
for all $n\in\mathbb{N}$, provided that $X$ and $Y$ are independent. In the rest of this paper, we may apply Lemma 2.20 without mentioning it.
In [16], Ruzsa studied some entropy inequalities that involve sums of random variables. In the following, more inequalities for discretized entropy are presented.
@@
Let $X,Y$ and $Z$ be three independent real random variables. Then, for all $n\in\mathbb{N}$,
@@
Without independence, the following lemma can be used as an alternative.
@@
Let $k,l\in\mathbb{Q}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there exist integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
@@
Let $k\in\mathbb{Z}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k)>0$ and $d_{2}=d_{2}(k)\geq 0$ such that
@@
For $k=0$, one can obtain by (8) that the conclusion holds for $d_{1}=d_{2}=1$. For $k=1$, the conclusion holds for $d_{1}=1$ and $d_{2}=0$. For k=1k=-1, Lemma 2.24 shows that we can take $d_{1}=3$ and $d_{2}=5$. Now we apply induction on $|k|$. Assume that the conclusion has been proved for $k$ with $|k|=m$, where $m\geq 1$. We consider the situation $k=\pm(m+1)$.
@@
Let $k,l\in\mathbb{Z}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}$ $\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\mathcal{I}_{n}(X,Y)+O_{k,l}(1).$
@@
When $k=l=0$, the choice $d_{1}=1$ and $d_{2}=0$ works. When $k\neq 0$ and $l=0$, the choice $d_{1}=d_{2}=1$ works. When $l\neq 0$, we apply Lemma 2.26 with the random variable $(X,lY)$ to get
@@
$\qquad\leq d_{1}(k,1)\left(H_{n}(X+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}(k,1)\mathcal{I}_{n}(X,Y)+O_{k,l}(1)$
for some $d_{1}(k,1)>0$ and $d_{2}(k,1)\geq 0$. Applying Lemma 2.26 again, with the random variable $(Y,X)$, we find that
@@
Since $k,l\in\mathbb{Q}$, we can write $k=k^{\prime}/t$ and $l=l^{\prime}/t$ for some $k^{\prime},l^{\prime}\in\mathbb{Z}$ and $t\in\mathbb{Z}^{\ast}$. It satisfies that $H_{n}(kX+lY)=H_{n}(k^{\prime}X+l^{\prime}Y)+O_{t}(1)$. The conclusion follows from Lemma 2.27. ∎
@@
for any $n\in\mathbb{N}$.
@@
Let $X,Y$ be independent random variables. Suppose that the real-valued functions $f,g$ are weakly $(\kappa,d)$-bi-Lipschitz on $range_{+}(X)$ and $range_{+}(Y)$, respectively. Suppose that $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$. Then
$H_{n}(f(X_{1})+f(X_{2})+g(Y_{1})+g(Y_{2}))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa)+O(\log d),$
$H_{n}(f(X_{1})-f(X_{2})+g(Y_{1})-g(Y_{2}))$ $\leq 7H_{n}(f(X)+g(Y))-3H_{n}(X)-3H_{n}(Y)+O(\log\kappa)+O(\log d),$
$H_{n}(f(X_{1})+f(X_{2})-g(Y_{1})-g(Y_{2}))$ $\leq 9H_{n}(f(X)+g(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\kappa)+O(\log d).$